# A new approach to delegate signing rights to proxy signers using isogeny-based cryptography

### v1

@@ Abstract
E-governance is a two-way protocol through which one can use government services, share data and request information. It refers to the use of communication and information technologies to provide government services to public in an efficient and fast manner. In addition, any document submitted to the e-Government system must be authenticated by a government officer using a digital signature scheme. In the context of digital signatures, the proxy signature is an important cryptographic primitive that allows the original signer to delegate signing authority to another signer (proxy signer). The proxy signature has a number of important applications in the e-government system. There are now a large amount of proxy signature schemes. The security of most of them relies on the following hard problems: the discrete logarithm problem and the factorization of integers problem. However, a large-scale quantum computer can solve them in polynomial time due to Shor’s algorithm. As a consequence, there is a need for a quantum computer-resistant proxy signature to secure e-governance system from quantum adversaries. In this work, we propose the first post-quantum isogeny based proxy signature scheme CSI-PS (commutative supersingular isogeny proxy signature). Our construction is proven to be uf-cma secure under the hardness of the group action inverse problem (GAIP) based on isogeny.
@@ 2.1 Elliptic curves and isogenies [17]
An isogeny $\chi:E_{1}\rightarrow E_{2}$ is an onto morphism inducing a group homomorphism from $E_{1}(\overline{\mathbb{F}}_{q})$ to $E_{2}(\overline{\mathbb{F}_{q}})$. Here $E_{1}$ and $E_{2}$ are called isogenous. Indeed, $E_{1}$ and $E_{2}$ having same number of points assure that there exist an isogeny between them. An isogeny sends the identity element $\Theta$ of $E_{1}$ to identity element $\Theta^{\prime}$ of $E_{2}$. An endomorphism is a morphism from $E$ to itself that fixes the distinguished point $\Theta$. The set of all endomorphisms on $E$ is denoted by $End(E)$. One should say that the set of all endomorphisms is actually a ring. The Frobenius endomorphism is an endomorphism on $E$ over $\mathbb{F}_{q}$ and it is defined as $\pi_{E}:(x:y:z)\longmapsto(x^{q}:y^{q}:z^{q})$. Another example of endomorphism is the multiplication-by-$n$ map, defined as $[n]:P\longmapsto[n]P$. Its kernel is called the $n$-th torsion subgroup $E[n]$.
The standard form of isogeny $\chi$ over $\mathbb{F}_{q}$ is given by $\chi(x,y)=\left(\frac{\alpha(x)}{\beta(x)},\frac{\gamma(x)}{\delta(x)}y\right)$, where $\alpha,\beta,\gamma,\delta\in\mathbb{F}_{q}[x]$, $\alpha$ is co-prime to $\beta$ and $\gamma$ is co-prime to $\delta$ in $\mathbb{F}_{q}[x]$. The degree of $\chi$ is denoted by $deg(\chi)$ and defined by $deg(\chi)=max\{deg(\alpha),deg(\beta)\}$. If $\left(\frac{\alpha(x)}{\beta(x)}\right)^{\prime}\neq 0$ then $\chi$ is called separable; otherwise, it is called inseparable. Note that the Frobenius endomorphism over $\mathbb{F}_{p}$ is an inseparable isogeny. The endomorphism $[m]$ for an elliptic curve $E/\mathbb{F}_{p}$ is separable if and only if $m$ and $p$ are co-prime. The $j$-invariant of $y^{2}=x^{3}+cx+d$ is defined by $j(E)=1728\frac{4c^{3}}{4c^{3}+27d^{2}}$. Two elliptic curves over $\overline{\mathbb{F}_{q}}$ with same $j$-invariant are called isomorphic. The torsion subgroup $E[p]$ is isomorphic to either $\mathbb{Z}/p\mathbb{Z}$ or 0. E is called ordinary in the first case, while $E$ is called a supersingular elliptic curve in the second case. If $E$ is supersingular over a finite field of characteristic $p$, then $j(E)$ lies in $\mathbb{F}_{p^{2}}$.
@@ 2.2 Endomorphism algebra and ideal class group [17]
A quaternion algebra over $\mathbb{F}_{q}$ is a $\mathbb{F}_{q}$-algebra with $\mathbb{F}_{q}$ basis $\{1,\alpha,\beta,\alpha\beta\}$, where $\alpha\beta=-\beta\alpha$ and $\alpha^{2},\beta^{2}\in\mathbb{F}_{q}-\{0\}$. Let $A$, $B$ be two rings and $f:A\rightarrow B$ be a ring homomorphism. An $A$-module structure on $B$ is defined by the scalar multiplication $a.b=f(a)b$. If $B$ has a ring structure along with an $A$-module structure via $f$ then $B$ is called an algebra. The endomorphism algebra of $E$ is denoted by $End^{0}(E):=End(E)\otimes_{\mathbb{Z}}\mathbb{Q}$, where $End^{0}(E)$ is a $\mathbb{Q}$-algebra. $End(E)$ is a free $\mathbb{Z}$-module of rank $r$. Note that $r=1$ implies that $End^{0}(E)$ isomorphic to $\mathbb{Q}$, $r=2$ implies that $End^{0}(E)$ is isomorphic to an imaginary quadratic field and $r=4$ implies that $End^{0}(E)$ is isomorphic to a quaternion algebra. Note that if $E$ is defined over a finite field, the case $r=1$ can not happen. An order $\vartheta$ of a $\mathbb{Q}$-algebra is a $\mathbb{Z}$-lattices of the $\mathbb{Q}$-algebra, which is also a subring. For any number field $K$, the ring of integers $\vartheta_{K}$ is the unique maximal order. The orders in imaginary quadratic number fields are defined by $\vartheta=\mathbb{Z}+f\vartheta_{K}$, where $f$ is an integer which is called the conductor of $\vartheta$.
The Frobenius endomorphism $\pi$ for an elliptic curve $E/\mathbb{F}_{q}$ satisfies the equation $\pi^{2}-t\pi+q=0$ with $|t|\leq 2\sqrt{q}$. Here $t$ is called trace of $\pi$ and $\Delta_{\pi}=t^{2}-4q$ is discriminant of $\pi$. Therefore, the discriminant $\Delta_{\pi}=t^{2}-4q$ of $\pi$ is negative and we have $\pi\in\mathbb{Q}(\sqrt{\Delta_{\pi}})$ which is an imaginary quadratic field.
Definition 2.
Let $G$ be a finite commutative group that acts freely and transitively on some set H. A hard homogeneous space (HHS) [20] consists of the sets $G$ and $H$, where the following computational problems are required to be easy:
  1. 1.
    Computation of the group action in $G$.
  2. 2.
    Random selection of an element from $G$ with uniform distribution.
  3. 3.
    Deciding validity and equality of a representation of H’s elements.
  4. 4.
    Computation of the group action of $g\in G$ on some $h\in H$.
and the following problems are assumed to be hard:
  1. 1.
    Given $h,k\in H$, finding $g\in G$ such that $k=gh$.
  2. 2.
    Given $h,k,w\in H$ with $k=gh$, finding $k^{\prime}=gw$
@@
Let $\vartheta$ be the order of a number field $K$. With the help of an $\vartheta$-ideal $L$, we can define a fractional $\vartheta$- ideal as $L^{\prime}=\omega L=\{\omega\alpha:\alpha\in L\}$ for some $\omega\in K^{*}$. A fractioal $\vartheta$-ideal $L^{\prime}$ is called invertible if there exists a fractional $\vartheta$-ideal $L^{\prime\prime}$ such that $L^{\prime}L^{\prime\prime}=\vartheta$. Collection of invertible fractional $\vartheta$-ideals forms a group with respect to multiplication of fractional $\vartheta$-ideals which is defined as $L^{\prime}N^{\prime}=\omega\omega^{\prime}LN$ for two fractional $\vartheta$-ideals $L^{\prime}=\omega L$ and $N^{\prime}=\omega^{\prime}N$. Let $I$ represents the group of invertible fractional $\vartheta$-ideals and $P$ denotes the subgroup of $I$ consisting of principal fractional $\vartheta$-ideals. The ideal class group is defined by the quotient group $cl(\vartheta)=\frac{I}{P}$. The $\mathbb{F}_{p}$-rational endomorphism ring $End_{p}(E)$ for a supersingular elliptic curve $E$ over $\mathbb{F}_{p}$ is an order in an imaginary quadratic field. Suppose $E_{p}$ is the set of isomorphism classes of elliptic curves over $\mathbb{F}_{p}$ whose $\mathbb{F}_{p}$-rational endomorphism ring is isomorphic to $\vartheta$. Then, we have a group action of $Cl(\vartheta)$ on $E_{p}$ which is free and transitive. As a consequence, $E_{p}$ is $Cl(\vartheta)$-torsor or principal homogeneous space (PHS).
@@
Let $E$ be a supersingular elliptic curve defined over $\mathbb{F}_{p}$ with the Frobenius endomorphism $\pi$, where $p\equiv 3\ (\textrm{mod}\ 4)$ and $p>3$. The necessary and sufficient condition for $End_{p}(E)=\mathbb{Z}[\pi]$ is that there exists a unique $e\in\mathbb{F}_{p}$ so that $E$ is $\mathbb{F}_{p}$ isomorphic to a Montgomery curve: $y^{2}=x^{3}+ex^{2}+x$, where $\mathbb{Z}[\pi]$ is a subring of $\mathbb{Q}(\sqrt{-p})$.
@@ 2.3 Commutative supersingular isogeny Diffie-Hellman (CSIDH) [23]
If $E$ is a supersingular curve over $\mathbb{F}_{p}$ then its ring of $\mathbb{F}_{p}$-rational endomorphism is an imaginary quadratic order $\vartheta\subseteq\mathbb{Q}(\sqrt{-p})$ which is denoted by $End_{p}(E)$. In CSIDH [23], we fix $E_{0}:y^{2}=x^{3}+x$ over $\mathbb{F}_{p}$ for the prime $p=4l_{1}\cdots l_{n}-1$, where $l_{1},\cdots,l_{n}$ are small distinct odd primes. As $p\equiv(-1)(\bmod~{}4)$, we can say that $E_{0}$ is supersingular. Hence $|E_{0}(\mathbb{F}_{p})|=p+1\equiv 0(\bmod~{}l_{i})$ and we can decompose $l_{i}\vartheta=J_{i}\overline{J_{i}}$, where $J_{i}=(l_{i},\pi-1)$ and $\overline{J_{i}}=(l_{i},\pi+1)$. Now we have to assume that $J_{i}$’s are evenly distributed in the class group and we expect that an ideal is of the form $J_{1}^{a_{1}}\cdots J_{n}^{a_{n}}$ with $(a_{1},\cdots a_{n})$ randomly chosen from $[-s,s]^{n}$ for some natural number $s$. Thereby, we may represent the ideal $J_{1}^{a_{1}}\cdots J_{n}^{a_{n}}$ as $(a_{1},\cdots,a_{n})$. It has been observed that it is sufficient to choose $m$ such that $2s+1\geq\sqrt[n]{|cl(\vartheta)}$. As the trace of Frobenius endomorphism is zero for supersingular elliptic curve over $\mathbb{F}_{p}$, its characteristic equation is $\pi^{2}+p=0$ yielding $\pi=\sqrt{-p}$. Thus, $End_{p}(E_{0})=\mathbb{Z}[\sqrt{-p}]=\mathbb{Z}[{\pi}]$. In other words, $E_{0}$ is $\mathbb{F}_{p}$ rational isomorphic to a Montgomery curve $E_{A}:y^{2}=x^{3}+ex^{2}+x$, which is represented by its coefficient $e$ by theorem 1.
@@ 2.6 SeaSign [26]
2.7 Notion of a proxy signature scheme [9]
@@ 4 Security Analysis
  • Unforgeability of proxy signature: We now prove that our proposed isogeny based proxy signature scheme CSI-PS is uf-cma secure under the hardness of the group action inverse problem of isogeny. $H_{1}$ and $H_{2}$ are considered as random oracles. We show this by contradiction. If possible let there be an adversary $\mathcal{A}$ with non-negligible advantage in the uf-cma game. Then we are going to show that using $\mathcal{A}$ an oracle machine $O^{\mathcal{A}}$ can be designed to break the isogeny problem by utilizing the outputs of $H_{1}$ and $H_{2}$. In order to do that, we present a series of games $Game^{0}$, $Game^{1}$, $Game^{2}$, $Game^{3}$, where $Game^{i}$ slightly modifies $Game^{i-1}$ for $i=1,2,3$. Let us denote the success probability of $\mathcal{A}$ in $Game^{i}$ as Prob[$Game^{i}$].

### v2

@@ Abstract
E-governance is a two-way protocol through which one can use government services, share data and request information. It refers to the use of communication and information technologies to provide government services to public in an efficient and fast manner. In addition, any document submitted to the e-Government system must be authenticated by a government officer using a digital signature scheme. In the context of digital signatures, the proxy signature is an important cryptographic primitive that allows the original signer to delegate signing authority to another signer (proxy signer). The proxy signature has a number of important applications in the e-government system. There are now a large amount of proxy signature schemes. The security of most of them relies on the following hard problems: the discrete logarithm problem and the factorization of integers problem. However, a large-scale quantum computers can solve them in polynomial time due to Shor’s algorithm. As a consequence, there is a need for a quantum computer-resistant proxy signature to secure e-governance system from quantum adversaries. In this work, we propose the first post-quantum isogeny based proxy signature scheme CSI-PS (commutative supersingular isogeny proxy signature). Our construction is proven to be uf-cma secure under the hardness of the group action inverse problem (GAIP) based on isogeny.
@@ 2.1 Elliptic curves and isogenies [17]
An isogeny $\chi:E_{1}\rightarrow E_{2}$ is an onto morphism inducing a group homomorphism from $E_{1}(\overline{\mathbb{F}}_{q})$ to $E_{2}(\overline{\mathbb{F}_{q}})$. Here $E_{1}$ and $E_{2}$ are called isogenous. Indeed, $E_{1}$ and $E_{2}$ having the same number of points assure an isogeny between them. An isogeny sends the identity element $\Theta$ of $E_{1}$ to identity element $\Theta^{\prime}$ of $E_{2}$. An endomorphism is a morphism from $E$ to itself that fixes the distinguished point $\Theta$. The set of all endomorphisms on $E$ is denoted by $\text{End}(E)$. One should say that the set of all endomorphisms is a ring. The Frobenius endomorphism is an endomorphism on $E$ over $\mathbb{F}_{q}$ and it is defined as $\pi_{E}:(x:y:z)\longmapsto(x^{q}:y^{q}:z^{q})$. Another example of an endomorphism is the multiplication-by-$n$ map, defined as $[n]:P\longmapsto[n]P$. Its kernel is called the $n$-th torsion subgroup $E[n]$.
The standard form of isogeny $\chi$ over $\mathbb{F}_{q}$ is given by $\chi(x,y)=\left(\frac{\alpha(x)}{\beta(x)},\frac{\gamma(x)}{\delta(x)}y\right)$, where $\alpha,\beta,\gamma,\delta\in\mathbb{F}_{q}[x]$, $\alpha$ is co-prime to $\beta$ and $\gamma$ is co-prime to $\delta$ in $\mathbb{F}_{q}[x]$. The degree of $\chi$ is denoted by $deg(\chi)$ and defined by $deg(\chi)=max\{deg(\alpha),deg(\beta)\}$. If $\left(\frac{\alpha(x)}{\beta(x)}\right)^{\prime}\neq 0$ then $\chi$ is called separable; otherwise, it is called inseparable. Note that the Frobenius endomorphism over $\mathbb{F}_{p}$ is an inseparable isogeny. The endomorphism $[m]$ for an elliptic curve $E/\mathbb{F}_{p}$ is separable if and only if $m$ and $p$ are co-prime. The $j$-invariant of $y^{2}=x^{3}+cx+d$ is defined by $j(E)=1728\frac{4c^{3}}{4c^{3}+27d^{2}}$. Two elliptic curves over $\overline{\mathbb{F}_{q}}$ with the same $j$-invariant are called isomorphic. The torsion subgroup $E[p]$ is isomorphic to either $\mathbb{Z}/p\mathbb{Z}$ or 0. E is called ordinary in the first case, while $E$ is called a supersingular elliptic curve in the second case. If $E$ is supersingular over a finite field of characteristic $p$, then $j(E)$ lies in $\mathbb{F}_{p^{2}}$.
@@ 2.2 Endomorphism algebra and ideal class group [17]
The endomorphism ring $\text{End}(E)$ is not just a $\mathbb{Z}$-module but also a $\mathbb{Z}$-algebra. The endomorphism algebra of $E$ is denoted by $\text{End}^{0}(E):=\text{End}(E)\otimes_{\mathbb{Z}}\mathbb{Q}$, making $\text{End}^{0}(E)$ a $\mathbb{Q}$-algebra. Since $\text{End}(E)$ is a free $\mathbb{Z}$-module of rank $r$, note the following cases: if $r=1$, then $\text{End}^{0}(E)$ is isomorphic to $\mathbb{Q}$; if $r=2$, then $\text{End}^{0}(E)$ is isomorphic to an imaginary quadratic field; and if $r=4$, then $\text{End}^{0}(E)$ is isomorphic to a quaternion algebra. If $E$ is defined over a finite field, the case $r=1$ cannot occur. An order $\vartheta$ of a $\mathbb{Q}$-algebra $K$ is a subring of $K$ that has the same rank as $K$. For any number field $K$, the ring of integers $\vartheta_{K}$ is the unique maximal order. In imaginary quadratic number fields, orders are defined by $\vartheta=\mathbb{Z}+f\vartheta_{K}$, where $f$ is an integer called the conductor of $\vartheta$.
@@
The Frobenius endomorphism $\pi$ for an elliptic curve $E/\mathbb{F}_{q}$ satisfies the equation $\pi^{2}-t\pi+q=0$ with $|t|\leq 2\sqrt{q}$. Here $t$ is called trace of $\pi$ and $\Delta_{\pi}=t^{2}-4q$ is discriminant of $\pi$. Therefore, the discriminant $\Delta_{\pi}=t^{2}-4q$ of $\pi$ is negative and we have $\pi\in\mathbb{Q}(\sqrt{\Delta_{\pi}})$ which is an imaginary quadratic field. Let $\vartheta$ be the order of a number field $K$. With the help of an $\vartheta$-ideal $L$, we can define a fractional $\vartheta$- ideal as $L^{\prime}=\omega L=\{\omega\alpha:\alpha\in L\}$ for some $\omega\in K^{*}$. A fractioal $\vartheta$-ideal $L^{\prime}$ is called invertible if there exists a fractional $\vartheta$-ideal $L^{\prime\prime}$ such that $L^{\prime}L^{\prime\prime}=\vartheta$. Collection of invertible fractional $\vartheta$-ideals forms a group with respect to multiplication of fractional $\vartheta$-ideals which is defined as $L^{\prime}N^{\prime}=\omega\omega^{\prime}LN$ for two fractional $\vartheta$-ideals $L^{\prime}=\omega L$ and $N^{\prime}=\omega^{\prime}N$. Let $I$ represent the group of invertible fractional $\vartheta$-ideals and $P$ denotes the subgroup of $I$ consisting of principal fractional $\vartheta$-ideals. The ideal class group is defined by the quotient group $cl(\vartheta)=\frac{I}{P}$. The $\mathbb{F}_{p}$-rational endomorphism ring $\text{End}_{p}(E)$ for a supersingular elliptic curve $E$ over $\mathbb{F}_{p}$ is an order in an imaginary quadratic field. Suppose $E_{p}$ is the set of isomorphism classes of elliptic curves over $\mathbb{F}_{p}$ whose $\mathbb{F}_{p}$-rational endomorphism ring is isomorphic to $\vartheta$. Then, we have a group action of $Cl(\vartheta)$ on $E_{p}$ which is free and transitive.
@@
Let $E$ be a supersingular elliptic curve defined over $\mathbb{F}_{p}$ with the Frobenius endomorphism $\pi$, where $p\equiv 3\ (\textrm{mod}\ 4)$ and $p>3$. The necessary and sufficient condition for $\text{End}_{p}(E)=\mathbb{Z}[\pi]$ is that there exists a unique $e\in\mathbb{F}_{p}$ so that $E$ is $\mathbb{F}_{p}$ isomorphic to a Montgomery curve: $y^{2}=x^{3}+ex^{2}+x$, where $\mathbb{Z}[\pi]$ is a subring of $\mathbb{Q}(\sqrt{-p})$.
@@ 2.3 Commutative supersingular isogeny Diffie-Hellman (CSIDH) [23]
If $E$ is a supersingular curve over $\mathbb{F}_{p}$ then its ring of $\mathbb{F}_{p}$-rational endomorphism is an imaginary quadratic order $\vartheta\subseteq\mathbb{Q}(\sqrt{-p})$ which is denoted by $\text{End}_{p}(E)$. In CSIDH [23], we fix $E_{0}:y^{2}=x^{3}+x$ over $\mathbb{F}_{p}$ for the prime $p=4l_{1}\cdots l_{n}-1$, where $l_{1},\cdots,l_{n}$ are small distinct odd primes. As $p\equiv(-1)(\bmod~{}4)$, we can say that $E_{0}$ is supersingular. Hence $|E_{0}(\mathbb{F}_{p})|=p+1\equiv 0(\bmod~{}l_{i})$ and we can decompose $l_{i}\vartheta=J_{i}\overline{J_{i}}$, where $J_{i}=(l_{i},\pi-1)$ and $\overline{J_{i}}=(l_{i},\pi+1)$. Now we have to assume that $J_{i}$’s are evenly distributed in the class group and we expect that an ideal is of the form $J_{1}^{a_{1}}\cdots J_{n}^{a_{n}}$ with $(a_{1},\cdots a_{n})$ randomly chosen from $[-s,s]^{n}$ for some natural number $s$. Thereby, we may represent the ideal $J_{1}^{a_{1}}\cdots J_{n}^{a_{n}}$ as $(a_{1},\cdots,a_{n})$. It has been observed that it is sufficient to choose $m$ such that $2s+1\geq\sqrt[n]{|cl(\vartheta)}$. As the trace of Frobenius endomorphism is zero for supersingular elliptic curve over $\mathbb{F}_{p}$, its characteristic equation is $\pi^{2}+p=0$ yielding $\pi=\sqrt{-p}$. Thus, $\text{End}_{p}(E_{0})=\mathbb{Z}[\sqrt{-p}]=\mathbb{Z}[{\pi}]$. In other words, $E_{0}$ is $\mathbb{F}_{p}$ rational isomorphic to a Montgomery curve $E_{A}:y^{2}=x^{3}+ex^{2}+x$, which is represented by its coefficient $e$ by theorem 1.
@@ 2.6 SeaSign [26]
2.7 Syntax of a proxy signature scheme [9]
@@ 4 Security analysis
  • Unforgeability of proxy signature: We now prove that our proposed isogeny based proxy signature scheme CSI-PS is uf-cma secure under the hardness of the group action inverse problem of isogeny. $H_{1}$ and $H_{2}$ are considered as random oracles. We show this by contradiction. If possible let there be an adversary $\mathcal{A}$ with a non-negligible advantage in the uf-cma game. Then we are going to show that using $\mathcal{A}$ an oracle machine $O^{\mathcal{A}}$ can be designed to break the isogeny problem by utilizing the outputs of $H_{1}$ and $H_{2}$. To do that, we present a series of games $Game^{0}$, $Game^{1}$, $Game^{2}$, $Game^{3}$, where $Game^{i}$ slightly modifies $Game^{i-1}$ for $i=1,2,3$. Let us denote the success probability of $\mathcal{A}$ in $Game^{i}$ as Prob[$Game^{i}$].